# Exposing a Fatal Flaw in Sample-based Quantum Diagonalization Methods

### v2

@@ V Conclusion
Alternatively, one may find systems such as the more strongly correlated iron-sulfur cluster [2Fe-2S], where sampling turns out to be less of a problem due to a more flat distribution of the CI coefficients. Unfortunately, the CI expansions generated by the QSCI method are much less compact than those generated by classical heuristics in these cases. HCI attains better energies with fewer determinants, leading to lower costs in the expensive diagonalization step compared to QSCI. This, again, limits the practical usefulness of QSCI.
@@ VI Acknowledgement
VII Data availability statement
@@
VIII Supporting Information Available
Effect of probability density function scaling (Eq. (4)), N2 ground-state energy error as a function of the number of quantum samples, and CI probability distribution plots for N2 and [2Fe-2S] systems. Examples of QSCI performance for 29 additional chemical systems.

### v3

@@ 5 Conclusion
Alternatively, one may find systems such as the more strongly correlated iron-sulfur cluster [2Fe-2S], where sampling turns out to be less of a problem due to a flatter distribution of the CI coefficients. Unfortunately, the CI expansions generated by the QSCI method are much less compact than those generated by classical heuristics in these cases. HCI attains better energies with fewer determinants, leading to lower costs in the expensive diagonalization step compared to QSCI. This, again, limits the practical usefulness of QSCI.
@@
Investigations into the system-size scaling show glimmers of hope for QSCI. Although both QSCI and HCI suffer from an exponential scaling with the system size, the exponent in the QSCI sampling routine appears to be lower than the cumulative determinant search and diagonalization costs for HCI. Unfortunately, to reach any practical advantage, one has to assume extremely optimistic improvements to quantum sampling rates.
For QSCI methods to have a practical advantage over classical SCI, clear gains must be demonstrated either in accuracy or efficiency. One possibility would be that QSCI could match or exceed the accuracy of classical SCI while keeping the cost of determinant selection low, giving better energies and wavefunctions without higher overhead. While we do find cases where the energy accuracy is comparable between QSCI and classical SCI, the sampling overheads for these cases are problematic. Alternatively, QSCI could offer value if its sampling procedure is markedly faster than classical methods at identifying important determinants, even if the final energies are somewhat less accurate. In principle, rapid sampling could outweigh modest accuracy losses. However, with current quantum devices operating at limited sampling rates, this potential advantage remains largely hypothetical. In regimes where classical SCI already yields accurate results at low cost, QSCI is unlikely to be competitive, regardless of quantum hardware improvements.
@@
Data availability statement
@@
{suppinfo}
Effect of probability density function scaling (Eq. (4)), N2 ground-state energy error as a function of the number of quantum samples, additional scaling factors for the N2 system, timings analysis QSCI and HCI, and CI probability distribution plots for N2 and [2Fe-2S] systems. Examples of QSCI performance for 29 additional chemical systems.